# Principal Component Copulas for Capital Modeling and Systemic Risk

### v2

@@ 4 Applications
We study the PCC and its estimators using simulated data and historical global market data. We start with a simulation experiment in high dimensions.
@@ 4.1 Simulation study
We perform a simulation study for a 100-dimensional PCC as schematically visualized in Fig. 1b. When simulating, the scheme is followed from right to left. We start with generating simulations for $\bm{P}$ based on the specified generator family. After this, we transform to $\bm{Y}$ using the linear transformation $\bm{Y}=\bm{W}\bm{P}$. Finally, we apply $F_{Y_{i}}$ to obtain samples for the copula $\bm{U}$. There are two different numerical approaches to obtain $F_{Y_{i}}$. The first method is to use Eqs. (15) and (17), which converges exponentially [16]. The second method is to simulate many observations for $Y_{i}$ and obtain an accurate simulation-based empirical distribution function $\hat{F}_{Y_{i}}$ for $Y_{i}$. Due to its fast convergence, we prefer and apply the first method.
The stochastic variables $Y_{i}$ have a mean equal to zero and a variance equal to $1$. Therefore, the first and second moments of $\bm{Y}$ are determined by the correlation matrix $\rho_{\bm{Y}}$. In our stylized example, the correlation matrix has the following structure:
@@
This parametrization leads to $d(d-1)/2$ different correlation parameters with values ranging from $0.10$ to $0.96$. By performing PCA, we obtain the true eigenvector matrix $\bm{W}$ and the true eigenvalues $\bm{\Lambda}$ of $\rho_{\bm{Y}}$. The first two PC vectors are shown in Fig. 2. These vectors are seen to describe a parallel shock and an anti-parallel shock. The first PCs have resulting eigenvalues $\lambda_{1}=43.8$ and $\lambda_{2}=18.5$, respectively, which therefore explain 62% of the variance. All higher PCs have eigenvalues less than 1, so that they do not carry relevant information.
@@
Here, ${\rm HB}(\alpha_{j},\beta_{j},\Lambda_{j})$ denotes the hyperbolic distribution from Eq. (3), where $\mu$ and $\delta$ are taken such that the mean is equal to zero and the variance is equal to $\Lambda_{j}$. In Table 1, the parameters selected for the study are presented. The chosen parameters for the first two PC distributions give rise to skewness and tail dependence, due to the heavier tails of the hyperbolic distribution compared to the normal distribution. The parameters lead to lower tail dependence estimates between risk pairs in the range from 0.05 to 0.85. We consider a higher dimension and a larger number of shape parameters in the simulations than in the case study of Section 4.2.3 in order to test the estimators under more challenging circumstances.
In the simulation study, we generate 1500 copula observations from the DGP for each estimation, which is a similar number of observations to our empirical case study in Section 4.2.3. As explained, we first sample the PCs independently from the hyperbolic and normal distributions with their true parameters. Then, we transform the simulated data using $\bm{Y}=\bm{W}\bm{P}$. Finally, we return $U_{i}=F_{Y_{i}}(Y_{i})$ for the simulated copula observations with $F_{Y_{i}}$ determined from Eqs. (12), (15) and (23). Having generated the copula data, we start with applying MLE for the 4 shape parameters $\alpha_{j}$ and $\beta_{j}$, taking the true eigenvector weights $w_{i,j}$ and eigenvalues $\Lambda_{j}$ as given. This hypothetical situation is used as performance reference. Next, the hybrid estimation procedure from Algorithm 1 is applied, which estimates all correlation parameters and the 4 shape parameters jointly. We do not inform our estimation algorithm about the underlying parameter structure.
In order to study the performance of the estimation procedure, 100 Monte Carlo replications of the full estimation procedure are generated and the mean and the standard deviation of the estimators are determined. In Table 1, the results for the four shape parameters using MLE and the results with Algorithm 1 are presented. We also show the results for the first two eigenvalues $\Lambda_{j}$. The estimation result for the corresponding eigenvectors $w_{i,j}$ are shown in Fig. 2, which reveals that the eigenvector estimates are very accurate. From Table 1, it is seen that the MLEs for the shape parameters perform well in terms of bias and standard errors. This is expected, since the true eigenvector weights and eigenvalues were used and since MLE is consistent with minimal asymptotic variance.
@@
$\lambda_{1}$ 43.78 43.74 (1.05)
$\lambda_{2}$ 18.53 18.59 (0.55)
$\alpha_{1}$ 0.50 0.51 (0.04) 0.53 (0.05)
$\beta_{1}$0.25 -0.26 (0.04) -0.27 (0.04)
$\alpha_{2}$ 1.00 0.94 (0.16) 0.95 (0.17)
$\beta_{2}$ 0.25 0.21 (0.09) 0.22 (0.10)
We compare these ML shape estimates with the hybrid GMM estimator of Algorithm 1, where we need less than five iterations for satisfactory convergence in this example. The estimation results of Algorithm 1 in the last columns are seen to give very limited performance loss compared to the ML shape estimates, even though in addition also all correlation parameters were estimated. We conclude that the algorithm works very well in high dimensions when the first PCs have large eigenvalues, so that these PC vectors can be robustly detected, avoiding the curse of dimensionality. Note that our example is stylized but realistic. In financial markets, we also typically observe a (few) dominant factor(s) in terms of variance. We have performed various simulation experiments with different distributions and different dimensionality, leading to similar results in terms of performance.
@@
In high-dimensions, it is possible to combine our algorithm with robust correlation estimators, such as eigenvalue clipping [21] or linear shrinkage [30]. In the first case, the eigenvalues of the lower PCs are replaced by a single eigenvalue such that the trace of the correlation matrix is preserved. This procedure lowers the number of parameters, because the correlation matrix is invariant for orthogonal transformations in the subspace of degenerate eigenvalues. Then, the resulting hyperbolic-normal copula from the example depends only on the parameters associated with the first two PCs, since the single degenerate eigenvalue for the lower PCs is fixed by $\Lambda_{>2}=d-\Lambda_{1}-\Lambda_{2}$.
@@ 4.2.1 Time series model
$r_{i,t}$ $=$ $\delta_{G,i}+\phi_{G,i}r_{i,t-1}+\sigma_{G,i,t}x_{i,t},\quad\sigma^{2}_{G,i,t}=\omega_{G,i}+\alpha_{G,i}x^{2}_{i,t-1}+\beta_{G,i}\sigma^{2}_{i,t-1}.$ (31)
@@ 4.2.2 Estimation results
We proceed with the return residuals $x_{i,t}$, which we assume to be i.i.d. The pseudo-copula observations are obtained by using the rank-based transform with $u_{i,t}=\text{rank}(x_{i,t})/(n+1)$ and $n=1303$ filtered return observations. We proceed with the inverse normal transform $\Phi^{-1}(u_{i,t})$ and apply PCA. The obtained eigenvectors and eigenvalues are used as starting point for Algorithm 1, which we apply to obtain the full iterative solution of the GMM estimator for the hyperbolic-normal (HB-N) PCC. The procedure from Algorithm 1 converges within a few iterations. The final vector weights are shown in Fig. 3. The variable indices have the same order as in Table 2.
The first PC-vector captures the movement of the market as a whole, since all vector weights $\hat{w}_{i,1}$ have the same sign. The Western European and the North American indices are found to have the strongest co-movement with the market, while the Maleysian and the Indonesian indices have the weakest co-movement. The second PC gives the most important mode of diversification. In our case study, it describes the movement of the Asian indices ($i=8$-14) compared to the Western indices ($i=1$-7, 15-16). The other indices ($i=17$-20) have small vector weights $\hat{w}_{i,2}$, signalling that they fall geographically or economically in between these two categories. We also note that the Japanese index ($i=9$) has a smaller vector weight $\hat{w}_{9,2}$ than the other Asian indices in the second PC, meaning that it moves less strongly opposite to the Western indices. We conclude that detailed global diversification effects are automatically captured by the copula. The third PC mainly describes the joint movement of the American indices ($i=15$-18) relative to other indices, where especially the additional variance of the two Latin-American indices ($i=17,18$) is captured.
@@
We perform the same estimation and simulation steps for the Skew $t_{1}$-$t_{d-1}$ PCC from Eqs. (9) and (11), leading to comparable visualisations of the copula. Finally, we introduce the Skew $t_{1}$-$t_{1}$ PCC, which is given by:
$P_{1}\sim{\rm Skew}_{1}\,t(\nu_{1},\Lambda_{1},\gamma),\quad P_{j}\sim t_{1}(\nu_{j},\Lambda_{j})\,\forall j\geq 2.$ (32)
Here, ${\rm Skew}\,t_{1}(\nu_{1},\Lambda_{1},\gamma_{1})$ and $t_{1}(\nu_{j},\Lambda_{j})$ denote one-dimensional distributions with degrees-of-freedom parameter $\nu_{j}$, asymmetry parameter $\gamma_{1}$, variance $\Lambda_{j}$ and zero mean. In case of the Skew $t_{1}$-$t_{1}$ PCC all generators are independent, where the first generator is modelled with the GH Skew $t_{1}$ distribution and the higher generators with the $t_{1}$ distribution. This PCC therefore has similarities to the Skew $t_{1}$-$t_{1}$ factor copula studied by Patton and Oh [8]. The stochastic representation for this copula can be obtained along the lines of Eq. (9) by allowing for an independent stochastic mixing variable $W_{j}$ for all higher generators $j$.
In Table 3, the estimation results for various 20-dimensional copulas are shown. In all cases, we use a single DoF parameter $\nu$ in such a way that the generators have a common tail coefficient. We consider the Gaussian copula, the $t_{d}$ copula, and the GH Skew $t_{d}$ copula [4] as benchmark copulas. The Gaussian copula and the $t_{d}$ copula are most commonly used in practice at financial institutions. The Skew $t_{d}$ copula is an important extension due to its ability to capture asymmetry between up and down movements [31]. It has also been a popular choice in recent dynamic factor copula applications due to its analytic tractability [11, 17, 19]. We estimate the Gaussian, the $t_{d}$, and the Skew $t_{d}$ copula with ML estimation [31]. The Skew $t_{d}$ copula depends on a vector $\bm{\gamma}$ of asymmetry parameters which we align with the first eigenvector of the normal-mixture covariance matrix, so that $\bm{\gamma}=\gamma\bm{w}_{1}$. This choice aligns with the other PCCs, and the parallel direction $\bm{w}_{1}$ is also the relevant direction for including skewness. To determine the standard deviations of estimated parameters, we reperform all estimation procedures for 100 bootstrap samples of filtered returns [32].
@@
The results for the Gaussian copula are shown in Table 3 as a reference for the other copulas. As expected, the AIC and BIC improve significantly for the $t_{d}$ copula, as it incorporates tail dependence, and for the Skew $t_{d}$ copula, as it incorporates asymmetric dependence [4, 17]. These normal mixture copulas are based on a single stochastic mixing variable. A novel aspect of our study is concerned with the optimal dimensionality of mixing variables. Recent factor copula studies have either assumed that all factors have one common mixing variable [9, 19, 11], or that all factors are independent [8]. The HB-N and the Skew $t_{1}$-$t_{1}$ PCCs are also based on independent generators. Patton and Oh have shown that such copula specifications lead to good performance for systemic risk [8], which we study in the next paragraph. However, these specifications are seen to be suboptimal in terms of likelihood. From Table 3, we see that the Skew $t_{1}$-$t_{d-1}$ PCC leads to significant improvements in AIC and BIC over the other copula specifications.
The interpretation for our results is that the Skew $t_{1}$-$t_{d-1}$ PCC benefits from having two stochastic mixing variables, namely $V$ for the first generator and $W$ for the higher generators in Eq. (9). The first generator drives the parallel movement of all market indices, as shown in Fig. 3, while the second and higher generators cause submarkets to move opposite to each other. The first generator is therefore important for systemic risk, while the second and higher generators lead to diversification. In crisis scenarios, a lack of diversification is expected. This can be captured by a high-dimensional copula where $V$ and $W$ are decoupled, so that they do not grow large simultaneously.
@@ 4.2.3 High-dimensional dependence and systemic risk
This measure is often used by financial institutions to assess pairwise dependence over a range of quantiles $q$. The limit $q\rightarrow 0$ of $\eta_{q,i,j}$ leads to the coefficient of lower tail dependence $\eta_{i,j}$. In Fig. 5, we show a pairwise CPJQE-plot for the FTSE-100 index and the S&P-500 index as a function of the quantile $q$. We compare the empirical estimates $\hat{\eta}_{q,i,j}$ based on historic pseudo-copula observations with simulated estimates $\tilde{\eta}_{q,i,j}$ based on various copula specifications. For the estimated copula models, $10^{6}$ simulations are used to accurately determine $\tilde{\eta}_{q,i,j}$. The confidence intervals for the empirical estimates are determined by using $1,000$ bootstrap samples of filtered returns. The Gaussian copula and the Student $t_{d}$ copula are seen to have lower pairwise dependence in the tail than the historic data. The Skew $t_{d}$, the HB-N PCC, the Skew $t_{1}$-$t_{d-1}$ PCC and the Skew $t_{1}$-$t_{1}$ PCC lead to an improved agreement with observations.
In order to assess systemic risk we are required to go beyond pairwise dependence. Instead of pairwize quantile exceedance, we consider the possibility of multiple market indices having a stressed downward return simultaneously. The fraction of market indices that have a joint downward exceedance of the quantile $q$ at time $t$ is given by the market distress ratio (MDR):
@@
When the MDR approaches one, it means that all market indices are simultaneously in distress beyond the quantile $q$. The probability that the MDR exceeds a prespecified ratio $k/d$ was introduced by Oh and Patton to study systemic risk in credit derivative markets [18]. A similar measure was used by Ouyang et al. to study systemic risk in commodity markets [19]. In both cases a dynamic copula was used to simulate the evolution of market distress probabilities over time.
Table 3: Estimatiom results for the copula shape parameters based on 20-dimensional weekly return data. Also the loglikelihood $\ell$, and the improvement in AIC and BIC over the Gaussian copula are presented. Bootstrap standard deviations of estimates are shown in parentheses. The best performing copula specification is highlighted in green.
Parameter Gauss Student $t_{d}$ Skew $t_{d}$ HB-N Skew $t_{1}$-$t_{1}$ Skew $t_{1}$-$t_{d-1}$
$\hat{\nu}$ 15.0  (1.2) 14.0   (1.3) 12.7   (1.4) 14.9 (1.1)
$\hat{\gamma}$2.74  (0.25) -2.16  (0.47) -2.84 (0.31)
$\hat{\alpha}$ 1.26   (0.32)
$\hat{\beta}$0.65   (0.22)
$\ell$ 10,505 10,895 11,001 10,668 10,812 11,041
$\Delta$ AIC -777 -988 -321 -610 -1068
$\Delta$ BIC -772 -978 -311 -599 -1058
@@
${\rm MDF}_{q,k,d}=\frac{1}{T}\sum_{t=1}^{T}{\rm MDI}_{t,q,k,d}=\frac{1}{T}\sum_{t=1}^{T}\bm{1}\left\{{\rm MDR}_{t,q,d}\geq\frac{k}{d}\right\}$ (35)
Here, the market distress indicator (MDI) specifies whether the stress event has occurred that $k$ or more market index returns simultaneously experienced a downward exceedance of the quantile $q$. The MDF is therefore an empirical counterpart of the joint distress probability considered by Oh and Patton [18].
@@
Table 4: Binomial test results ($p$-values) for historical counts of the market distress indicator ${\rm MDI}_{t,q,k,d}$ based on 1303 20-dimensional filtered weekly return observations for three commonly used high-dimensonal copulas and three PCCs. The $p$-values below and above a threshold of 5% are coloured red and green.
Specification Count Gauss $t_{d}$ Skew $t_{d}$ HB-N Skew $t_{1}$-$t_{1}$ Skew $t_{1}$-$t_{d-1}$
$q=0.15$, $k/d=1$ 9 $<10^{-4}$ 0.0006 0.014 0.16 0.27 0.41
$q=0.20$, $k/d=1$ 16 $<10^{-4}$ $<10^{-4}$ 0.002 0.06 0.12 0.23
$q=0.15$, $k/d=0.9$ 28 $<10^{-4}$ 0.0002 0.013 0.14 0.19 0.31
$q=0.20$, $k/d=0.9$ 44 $<10^{-4}$ 0.0001 0.009 0.11 0.13 0.26

### v3

@@ 4 Applications
We study the PCC and its estimators using simulated data and historical global market data.
@@ 4.1 Simulation study
We perform a simulation study for a 100-dimensional PCC as schematically visualized in Fig. 3b. When simulating, the scheme is followed from right to left. We start with generating simulations for $\bm{P}$ based on the specified generator family. After this, we transform to $\bm{Y}$ using the linear transformation $\bm{Y}=\bm{W}\bm{P}$. Finally, we apply $F_{Y_{i}}$ to obtain samples for the copula $\bm{U}$. The stochastic variables $Y_{i}$ have a mean equal to zero and a variance equal to $1$. Therefore, the first and second moments of $\bm{Y}$ are determined by the correlation matrix $\rho_{\bm{Y}}$. In our stylized example, the correlation matrix has the following structure:
@@
This parametrization leads to $d(d-1)/2$ different correlation parameters with values ranging from $0.10$ to $0.96$. By performing PCA, we obtain the true eigenvector matrix $\bm{W}$ and the true eigenvalues $\bm{\Lambda}$ of $\rho_{\bm{Y}}$. The first two PC vectors are shown in Fig. 4. These vectors are seen to describe a parallel shock and an anti-parallel shock. The first PCs have resulting eigenvalues $\lambda_{1}=43.6$ and $\lambda_{2}=18.7$, respectively, which therefore explain 62% of the variance. All higher PCs have eigenvalues less than 1, so that they do not carry relevant information.
@@
Here, ${\rm HB}(\alpha_{j},\beta_{j},\Lambda_{j})$ denotes the hyperbolic distribution from Eq. (3), where $\mu$ and $\chi$ are taken such that the mean is equal to zero and the variance is equal to $\Lambda_{j}$. In Table 1, the parameters selected for the study are presented. The chosen parameters for the first two PC distributions give rise to skewness and tail dependence, due to the heavier tails of the hyperbolic distribution compared to the normal distribution. The parameters lead to lower tail dependence estimates between risk pairs in the range from 0.05 to 0.85. We consider a higher dimension and a larger number of shape parameters in the simulations than in the case study of Section 4.2 in order to test the estimators under more challenging circumstances.
In the simulation study, we generate 1500 copula observations for each estimation, which is a similar number of observations to our empirical case study in Section 4.2. Following the specified DGP, we sample the PCs $\bm{P}$ independently from the hyperbolic and normal distributions with their true parameters $\alpha_{j}$, $\beta_{j}$ and $\Lambda_{j}$. Then, we transform the simulated data using $\bm{Y}=\bm{W}\bm{P}$ with $\bm{W}$ the true eigenvector matrix of $\rho_{\bm{Y}}$. Finally, we return $U_{i}=F_{Y_{i}}(Y_{i})$ for the simulated copula observations. There are two different numerical approaches to obtain $F_{Y_{i}}$. The first method is to use Eqs. (14) and (22), which converges exponentially [16]. The second method is to simulate many observations for $Y_{i}$ and obtain a simulation-based empirical distribution function $\hat{F}_{Y_{i}}$ for $Y_{i}$. Due to its fast convergence, we use the first method. Having generated the copula data, we start with applying MLE for the 4 shape parameters $\hat{\alpha}_{j}$ and $\hat{\beta}_{j}$, taking the true eigenvector weights $w_{i,j}$ and eigenvalues $\Lambda_{j}$ as given. This hypothetical situation is used as performance reference. Next, the hybrid estimation procedure from Algorithm 1 is applied, which estimates all correlation parameters and the 4 shape parameters jointly. We do not inform our estimation algorithm about the underlying parameter structure.
In order to study the performance of the estimation procedure, 100 Monte Carlo replications of the full estimation procedure are generated and the mean and the standard deviation of the estimators are determined. In Table 1, the results for the four shape parameters using MLE and the results with Algorithm 1 are presented. We also show the results for the first two eigenvalues $\hat{\Lambda}_{j}$. The estimation result for the corresponding eigenvectors $\hat{w}_{i,j}$ are shown in Fig. 4, which reveals that the eigenvector estimates are very accurate. From Table 1, it is seen that the MLEs for the shape parameters perform well in terms of bias and standard errors. This is expected, since the true eigenvector weights and eigenvalues were used and since MLE is consistent with minimal asymptotic variance.
@@
$\lambda_{1}$ 43.61 43.57 (1.07)
$\lambda_{2}$ 18.70 18.77 (0.55)
$\alpha_{1}$ 0.50 0.47 (0.05) 0.47 (0.06)
$\beta_{1}$0.25 -0.23 (0.04) -0.23 (0.05)
$\alpha_{2}$ 1.00 0.99 (0.04) 1.02 (0.07)
$\beta_{2}$ 0.25 0.26 (0.06) 0.26 (0.07)
We compare these ML shape estimates with the hybrid GMM estimator of Algorithm 1, where we need less than five iterations for satisfactory convergence in this example. The estimation results of Algorithm 1 in the last columns are seen to give limited performance loss compared to the ML shape estimates, even though in addition also all correlation parameters were estimated. We conclude that the algorithm works well in high dimensions when the first PCs have large eigenvalues, so that these PC vectors can be robustly detected, avoiding the curse of dimensionality. Note that our example is stylized but realistic. In financial markets, we also typically observe a (few) dominant factor(s) in terms of variance. We have performed various simulation experiments with different distributions and different dimensionality, leading to similar results in terms of performance.
@@
In high-dimensions, it is possible to combine our algorithm with robust correlation estimators, such as eigenvalue clipping [21] or shrinkage [29]. In the first case, the eigenvalues of the lower PCs are replaced by a single eigenvalue such that the trace of the correlation matrix is preserved. This procedure lowers the number of parameters, because the correlation matrix is invariant for orthogonal transformations in the subspace of degenerate eigenvalues. Adjusting the eigenvalues with clipping or shrinkage is advised when the number of dimensions approaches the number of data points $d/n>0.1$. [29]
@@ 4.2.1 Time series model
$r_{i,t}$ $=$ $\delta_{G,i}+\phi_{G,i}r_{i,t-1}+\epsilon_{i,t},\quad\sigma^{2}_{G,i,t}=\omega_{G,i}+\alpha_{G,i}\epsilon^{2}_{i,t-1}+\beta_{G,i}\sigma^{2}_{i,t-1}.$
@@ 4.2.2 Estimation results
We proceed with the devolatized residuals $x_{i,t}=\epsilon_{i,t}/\sigma_{G,i,t}$, which we assume to be i.i.d. The pseudo-copula observations are obtained by using the rank-based transform with $u_{i,t}=\text{rank}(x_{i,t})/(n+1)$ and $n=1303$ filtered return observations. We initialize the estimation procedure using the inverse normal transform $\Phi^{-1}(u_{i,t})$ and applying PCA. The obtained eigenvectors and eigenvalues are used as starting point for Algorithm 1, which we apply to obtain the full iterative solution of the GMM estimator for the hyperbolic-normal (HB-N) PCC. The procedure from Algorithm 1 converges within a few iterations. The final vector weights are shown in Fig. 5. The variable indices have the same order as in Table 2.
The first PC-vector captures the movement of the market as a whole, since all vector weights $\hat{w}_{i,1}$ have the same sign. The Western European and the North American indices are found to have the strongest co-movement, while the Maleysian and the Indonesian indices have the weakest co-movement. The second PC gives the most important mode of diversification. In our case study, it describes the movement of the Asian indices ($i=8$-14) compared to the Western indices ($i=1$-7, 15-16). The other indices ($i=17$-20) have small vector weights $\hat{w}_{i,2}$, signalling that they fall geographically or economically in between these two categories. We also note that the Japanese index ($i=9$) has a smaller vector weight $\hat{w}_{9,2}$ than the other Asian indices in the second PC, meaning that it moves less strongly opposite to the Western indices. We conclude that detailed global diversification effects are automatically captured by the copula. The third PC mainly describes the joint movement of the American indices ($i=15$-18) relative to other indices, where especially the additional variance of the two Latin-American indices ($i=17,18$) is captured.
@@
We perform the same estimation and simulation steps also for the Skew $t_{1}$-$t_{d-1}$ PCC from Eqs. (9) and (11), leading to comparable visualisations of the copula. Finally, we introduce the Skew $t_{1}$-$t_{1}$ PCC, which is given by:
$P_{1}\sim{\rm Skew}\,t_{1}(\nu_{1},\Lambda_{1},\gamma_{1}),\quad P_{j}\sim t_{1}(\nu_{j},\Lambda_{j})\,\forall j\geq 2.$ (30)
Here, ${\rm Skew}\,t_{1}(\nu_{1},\Lambda_{1},\gamma_{1})$ and $t_{1}(\nu_{j},\Lambda_{j})$ denote one-dimensional distributions with degrees-of-freedom parameter $\nu_{j}$, asymmetry parameter $\gamma_{1}$, variance $\Lambda_{j}$ and zero mean. In case of the Skew $t_{1}$-$t_{1}$ PCC all generators are independent, where the first generator is modelled with the GH skew $t_{1}$ distribution and the higher generators with the $t_{1}$ distribution. This PCC therefore has similarities to the Skew $t_{1}$-$t_{1}$ factor copula studied by [8]. The stochastic representation for this copula can be obtained along the lines of Eq. (9) using an independent stochastic mixing variable $W_{j}$ for each of the higher generators $j$.
@@
In Table 3, the estimation results for various 20-dimensional copulas are shown. In all cases, we use a single DoF parameter $\nu$ in such a way that the generators have a common tail coefficient. We consider the Gaussian copula, the $t_{d}$ copula, and the GH skew $t_{d}$ copula as benchmark copulas. The Gaussian copula and the $t_{d}$ copula are most commonly used in practice at financial institutions. The Skew $t_{d}$ copula is an important extension due to its ability to capture asymmetry between up and down movements [30]. It has also been a popular choice in recent dynamic factor copula applications due to its analytic tractability [17, 19, 11]. We estimate the Gaussian, the $t_{d}$, and the Skew $t_{d}$ copula with ML estimation [30]. The Skew $t_{d}$ copula depends on a vector $\bm{\gamma}$ of asymmetry parameters which we align with the first eigenvector of the normal-mixture covariance matrix, so that $\bm{\gamma}=\gamma\bm{w}_{1}$. This choice aligns with the other PCCs, and the parallel direction $\bm{w}_{1}$ is also the relevant direction for including skewness in our case study. To determine the standard deviations of estimated parameters, we reperform all estimation procedures for 100 bootstrap samples of filtered returns [31].
The results for the Gaussian copula are shown in Table 3 as a reference for the other copulas. As expected, the AIC and BIC improve significantly for the $t_{d}$ copula, as it incorporates tail dependence, and for the Skew $t_{d}$ copula, as it incorporates asymmetric dependence [17]. These normal mixture copulas are based on a single stochastic mixing variable. A novel aspect of our study is concerned with the optimal dimensionality of mixing variables. Recent factor copula studies have either assumed that all factors have one common mixing variable [9, 19, 11], or that all factors are independent [8]. The HB-N and the Skew $t_{1}$-$t_{1}$ PCCs are also based on independent generators. [8] have shown that such copula specifications lead to good performance for systemic risk, which we study in the next paragraph. However, these specifications are seen to be suboptimal in terms of likelihood. From Table 3, we see that the Skew $t_{1}$-$t_{d-1}$ PCC leads to significant improvements in AIC and BIC over the other copula specifications.
The interpretation for our results is that the Skew $t_{1}$-$t_{d-1}$ PCC benefits from having two stochastic mixing variables, namely $V$ for the first generator and $W$ for the higher generators in Eq. (9). The first generator drives the parallel movement of all market indices, as shown in Fig. 5, while the second and higher generators cause submarkets to move opposite to each other. The first generator is therefore important for systemic risk, while the second and higher generators lead to diversification. In crisis scenarios, a lack of diversification is expected, where all submarkets move in the same direction. This can be captured by a high-dimensional copula where $V$ and $W$ are decoupled, so that they do not grow large simultaneously.
Table 3: Estimation results for the copula shape parameters based on 20-dimensional weekly return data. Also the loglikelihood $\ell$, and the improvement in AIC and BIC over the Gaussian copula are presented. Bootstrap standard deviations of estimates are shown in parentheses. The best performing copula, the Skew $t_{1}$-$t_{d-1}$ PCC, is highlighted in bold.
Gauss $t_{d}$ Skew $t_{d}$ HB-N Skew $t_{1}$-$t_{1}$ Skew $t_{1}$-$t_{d-1}$
$\hat{\nu}$ 15.1 (1.2) 14.0 (1.3) 12.7 (1.4) 14.9 (1.1)
$\hat{\gamma}$2.7 (0.2) -2.2 (0.5) -2.8 (0.3)
$\hat{\alpha}$ 1.3 (0.3)
$\hat{\beta}$0.7 (0.2)
$\ell$ 10,505 10,895 11,001 10,668 10,812 11,041
$\Delta$ AIC -777 -988 -321 -610 -1068
$\Delta$ BIC -772 -978 -311 -599 -1058
@@ 4.2.3 High-dimensional dependence and systemic risk
This measure is often used by financial institutions to assess pairwise dependence over a range of quantiles $q$. The limit $q\rightarrow 0$ of $\eta_{q,i,j}$ leads to the coefficient of lower tail dependence $\eta_{i,j}$. In Fig. 7, we show a pairwise CPJQE-plot for the FTSE-100 index and the S&P-500 index as a function of the quantile $q$. We compare the empirical estimates $\hat{\eta}_{q,i,j}$ based on historic pseudo-copula observations with simulated estimates $\tilde{\eta}_{q,i,j}$ based on various copula specifications. For the estimated copula models, $10^{6}$ simulations are used to accurately determine $\tilde{\eta}_{q,i,j}$. The confidence intervals for the empirical estimates are determined by using $1,000$ bootstrap samples of filtered returns. The Gaussian copula and the Student $t_{d}$ copula are seen to have lower pairwise dependence in the tail than the historic data. The Skew $t_{d}$, the HB-N PCC, the Skew $t_{1}$-$t_{d-1}$ PCC and the Skew $t_{1}$-$t_{1}$ PCC lead to improved agreement.
In order to assess systemic risk we are required to go beyond pairwise dependence. Instead of pairwise quantile exceedance, we consider the possibility of multiple market indices having a stressed downward return simultaneously. The fraction of market indices that have a joint downward exceedance of the quantile $q$ at time $t$ is given by the market distress ratio (MDR):
@@
When the MDR approaches one, it means that all market indices are simultaneously in distress beyond the quantile $q$. The probability that the MDR exceeds a prespecified ratio $k/d$ was introduced by [18] to study systemic risk in credit derivative markets. A similar measure was used by [19] to study systemic risk in commodity markets. In both cases a dynamic copula was used to simulate the evolution of market distress probabilities over time.
@@
${\rm MDF}_{q,k,d}=\frac{1}{n}\sum_{t=1}^{n}{\rm MDI}_{t,q,k,d}=\frac{1}{n}\sum_{t=1}^{n}\bm{1}\left\{{\rm MDR}_{t,q,d}\geq\frac{k}{d}\right\}$ (33)
Here, the market distress indicator (MDI) specifies whether the stress event has occurred that $k$ or more market index returns simultaneously experienced a downward exceedance of the quantile $q$. The MDF is therefore an empirical counterpart of the joint distress probability considered by [18].
@@
Table 4: Binomial test results ($p$-values) for historical counts (##\##) of the market distress indicator ${\rm MDI}_{t,q,k,d}$ based on 1303 filtered weekly observations of 20 major stock indices using three common high-dimensonal copulas and three PCCs. The test is performed for $d=20$, and various combinations of $q$ and $k/d$. The normal mixture copulas would be rejected, while the PCCs would be accepted at a significance level of 5%.
$(q,k/d)$#\## Gauss $t_{d}$ Skew $t_{d}$ HB-N Skew $t_{1}$-$t_{1}$ Skew $t_{1}$-$t_{d-1}$
(0.15,1) 9 $<10^{-4}$ 0.0006 0.014 0.16 0.27 0.41
(0.2,1) 16 $<10^{-4}$ $<10^{-4}$ 0.002 0.06 0.12 0.23
(0.15,0.9) 28 $<10^{-4}$ 0.0002 0.013 0.14 0.19 0.31
(0.2,0.9) 44 $<10^{-4}$ 0.0001 0.009 0.11 0.13 0.26